OpenStudy (anonymous):

Which algebraic property could be used to rewrite 3x ⋅ (7y ⋅ 4) as (3x ⋅ 7y) ⋅ 4? Associative Property of Addition Associative Property of Multiplication Commutative Property of Addition Commutative Property of Multiplication

OpenStudy (anonymous):

Medal and fan :) @Nnesha @mathway @RosieF @helper1212

Mehek (mehek14):

since there is multiplication, the answer can't be A or C

Mehek (mehek14):

That leaves us with B and D

Mehek (mehek14):

Associative Property of Multiplication 

Mehek (mehek14):

Commutative Property of Multiplication

OpenStudy (anonymous):

D?

Mehek (mehek14):



Mehek (mehek14):

\(\bf{Associative ~Property ~of ~Multiplication \rightarrow {a*(b*c)=(a*b)*c}}\\\bf{Commutative~ Property~ of~ Multiplication\rightarrow{a*b=b*a}} \)